# El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data

### v1

@@ 5.4 Lensed image families and clumps in the source
Figure 11: Multiple images of peculiar features in the background source galaxy. Top row: families of multiple images in image plane together with the critical lines and the observed flux from the arc (after lens subtraction in filters F [^4] W ad F200W). Also shown are the positions of the three deflectors with pink crosses (residuals remaining after lens light subtraction are visible around L1 and L2 positions), as well as the critical lines with dashed white lines. Our model also predicts images at the position of L1 and L2, although there are not shown here to limit clutter. Bottom row: corresponding image positions on the source plane, superimposed onto one sample of our source model. The tangential caustic is shown with dotted white lines. Note that each column has its own dynamic range.
@@
We are able to confirm that parts of the background galaxy are multiply imaged five times (excluding central images); in Fig. 11 this is shown by image families 1 and 3. This peculiar lensing configuration has been first suspected in Kamieneski et al. (2023). We check if these fifth images are magnified by computing their model-predicted magnification from our fiducial mass model. We obtain magnifications of $\mu\approx-3$ for image 1d and $\mu\approx-5$ for image 3b, thus we find there are magnified. We note that this “exotic” lensing configuration typically arises in systems with multiple deflectors and has been the topic of several theoretical works in the past (e.g., Shin & Evans 2008; Orban de Xivry & Marshall 2009; Bozza et al. 2020). We briefly discuss in Appendix A.4 potential implications regarding this lensing configuration.
@@
We find that one of the images used in Kamieneski et al. (2023) as being part of the image family 1 (their image 1b) in fact does not belong to this set of multiple image according to our lens model. On the one hand, we find that their image 1d is in fact composed of the two images denoted by 1a and 1c, which appear like merging while crossing the tangential critical line. On the other hand, their image 1b corresponds to our image 4a in Fig. 11, and thus originates from a distinct compact bright clump in the source, located at the north of the central bulge. We note that this clump has similar colors as the bulge of the source galaxy, which is why it could have been incorrectly attributed to a lensed image of the bulge. Interestingly, not all images of this clump are visible in the data. We believe it is mostly due to a combination of a the larger FWHM of the PSF in redder bands, and stronger contamination by the deflectors light. To better visualize this, we show in Fig. 12 a comparison between our the lensed and convolved source model and the corresponding model before convolution by the PSF (in F356W although it is similar in other filters). We see that only image 4a remains clearly visible after convolution. Moreover, the lensed images appear either close to the bright bulge of the source (image 4d), or close to regions heavily contaminated by the deflectors light (images 4b and 4c); therefore, they are not seen in the data. From one of our a higher resolution source models (with $n_{\rm src}=300$), we estimate that the physical size of the clump is approximately $400$ pc at redshift $z_{\rm s}=2.291$. While this estimate is an upper-bound to the true size—since increasing the source resolution tends to make it more compact—this clump in the source is remarkably small. It is comparable in size to some of the structures in the jets of lensed sources analyzed in Powell et al. (2022) using VLBI observations with milli-arcsecond resolution (against 30 milli-arcsecond resolution with JWST/NIRCam data).
@@
Finally, we note that our extended source model can be used to extract the spectral energy distribution (SED) of compact unresolved structures in the arc. We show in Appendix B the case of one of the brightest point-like feature in El Anzuelo (image 2b in Fig. 11). As our pixelated source model, at its fiducial resolution ($n_{\rm src}=100$), does not capture unresolved features by construction, it can be directly subtracted from the data to separate the extended source component from these unresolved features. It is then easier to perform photometric measurements and estimate the SED of such compact source regions, which are important tracers of the star formation history (as shown in Kamieneski et al. 2023).
@@ 5.6 Release of lens models within the COOLEST standard
Previous sections describe our modeling and inference approach with a level of details that aim at maximizing the reproducibility of our analysis. With this idea in mind, we also publicly release all lens models [^13] Lens models under the COOLEST format will be released at the publication of the article, although they can be shared earlier upon reasonable request. following the strong lensing standard COOLEST (Galan et al. 2023). These models can then be easily loaded using the Python package coolest [^14] https://github.com/aymgal/COOLEST for further analyses, in particular with any modeling codes that share an interface with COOLEST (see the documentation for a list). We encourage future lens models to be released under this standard, in order to facilitate comparison with the present analysis.
@@ 6.1 Mass density profile of elliptical galaxies
Gravitational lensing allows us to understand the interplay between dark matter and baryons, and how this interplay shapes the inner structure of galaxies and galaxy clusters. Strong lensing in particular probes the two-dimensional mass distribution of galaxies via the enclosed mass within the Einstein radius $\theta_{\rm E}$ and—when high S/N extended arcs are used—the radial density slope measured at $\theta_{\rm E}$. Different processes can impact the density slope, such as the cooling of baryonic matter followed by dark matter halo contraction (e.g., Gnedin et al. 2004; Abadi et al. 2010), itself compensated for by outflows from stellar winds or AGN feedback (e.g., Johansson et al. 2012; Dubois et al. 2013). Large elliptical galaxies, which are prevalent of cluster-scale and galaxy-scale strong lenses, were found to have a radial behavior remarkably close to an isothermal profile (i.e., $\gamma=2$) from lensing and lens stellar dynamics constraints (e.g., Treu & Koopmans 2002; Auger et al. 2010; Sonnenfeld et al. 2013). This result has been termed the “bulge-halo conspiracy” in the literature (Treu et al. 2006; Dutton & Treu 2014; Etherington et al. 2023a): despite the different radial behaviors of baryonic and dark matter profiles, their combination seems to lead to an isothermal density profile, with no evident signs for evolution with redshift (e.g., Sonnenfeld et al. 2013; Etherington et al. 2023a). At the population level, this result has been recently confirmed using lensing-only constraints, larger samples, and more modern lens modeling techniques (Shajib et al. 2021; Etherington et al. 2022; Tan et al. 2023).
The three foreground deflectors if El Anzuelo have typical photometric signatures of elliptical galaxies, and have been confirmed to be members of the massive cluster El Gordo (e.g., Caminha et al. 2023). In this work, we model the mass distribution of these galaxies using elliptical power-law profiles perturbed by the cluster mass distribution and an additional external shear. As presented in Sect. 5 (Table 2), we infer logarithmic density slopes for L1 and L2 of $\gamma_{\rm L1}=2.23\pm 0.05$ and $\gamma_{\rm L2}=2.21\pm 0.04$, respectively. These values are inconsistent with an isothermal slope, and $\gamma_{\rm L2}$ is $2.7\sigma$ away from the mean value $\langle\gamma\rangle=2.075\pm 0.024$ of Etherington et al. (2023a) inferred from a subset of SLACS (Bolton et al. 2006) and GALLERY (Shu et al. 2016) systems. As detailed in Sect. 4, we infer our uncertainties by marginalizing over an ensemble of model variations with Bayesian weighting. We also looked for potential systematic effects and did not find statistically significant changes in the posteriors distributions by running additional models. Here, we first compare the precision of these measurements with previous analyses of similar imaging data, then discuss the possible origins of these results in terms of galaxy structure.
@@
That elliptical galaxies have nearly isothermal radial density profiles in their inner region has been observed in isolated deflectors, such as those from the SLACS (Bolton et al. 2006) and SL2S (Gavazzi et al. 2012) samples. The three main deflectors of El Anzuelo are, however, not isolated but genuine members of El Gordo that are gravitationally interacting. Therefore, it may not be surprising to measure density slopes that deviate from an isothermal profile. For most of isolated lens systems from SLACS (or comparable samples) we note that the ratio $\bar{r}_{\rm E/eff}$ between the Einstein radius and the half-light radius, $\bar{r}_{\rm E/eff}\equiv\theta_{\rm E}/\theta_{\rm eff}\lesssim 1$, meaning that lensing overall constrains the inner region of the lens galaxies, within their half-light radius (e.g., see the middle panel of Fig. 15 from Etherington et al. 2022, for SLACS and GALLERY samples). For the case of El Anzuelo, L1 and L2 both have $\bar{r}_{\rm E/eff,L1}\approx 1.6$ and $\bar{r}_{\rm E/eff,L2}\approx 1.3$ respectively (as visualized in Fig. 10), meaning that lensing provides constraints at overall larger radii compared to lens galaxies from the aforementioned samples [^16] The effective radius for L2 may be underestimated due to due blending with the lensed arcs (see Appendix A.1). However, due to the lensing boost by the cluster, the ratio $\bar{r}_{\rm E/eff,\rm L2}$ is likely larger than unity.. The main reason for El Anzuelo deflectors to have $\bar{r}>1$ is that their apparent lensing effect is boosted by the mass density of the host cluster, leading to the appearance of an Einstein ring that is larger than if L1 or L2 were isolated galaxies. This effect can be qualitatively confirmed by noting the difference in $\theta_{\rm E}$ values between the “Shear only” and “Cluster+Shear” models (e.g., Fig. 9): including the cluster mass distribution in the model significantly reduces $\theta_{\rm E}$ values for L1 and L2 in order to fit the imaging data. Therefore, the observed lensing features probe the mass distribution of the deflectors at sensibly larger radii, compared to isolated deflectors.
@@
6.2 Can the host cluster be approximated by a constant external shear?
Figure 13: Shear orientation and strength over the field of view of El Anzuelo. For the “Shear only” model, the shear field is a uniform external shear with strength $\gamma_{\rm ext}$; for the “Cluster only”, it is the shear field from the cluster model of Caminha et al. (2023); for the “Cluster+Shear” model, it is the combined effect from the uniform external shear and the cluster shear. For the latter two models, the averaged shear strength $\langle\gamma_{\rm shear}\rangle$ is also indicated the legend. The absolute length of the shear markers is arbitrary, but their relative size is respected. The blue ellipse indicates the orientation of the mass density of L1. On the left of the plot, we show the directions pointing towards the two main dark matter halo component of the El Gordo cluster. Note that the shear orientation is consistent with the position of the cluster scale components.
@@ 6.3 Using extended arcs instead of point-like constraints
We compared in more details our results with the lens model of Kamieneski et al. (2023), especially for three image families. In their analysis, the authors used the 8 NIRCam filters to look for compact features in image plane that have consistent colors, and considered those as multiple images of same region in the source. For their image family 1, they used a simplified initial lens model to confirm the five-image configuration, interpreted as a confirmation of the origin of images 1a to 1e being lensed versions of the bright central region of the source galaxy. In Sect. 5.4, we show that image 1b does not originate from the center of the source, but rather from an off-centered compact clump of size $\lesssim 400$ pc. In NIRCam images, this clump has similar colors as the central bulge of the source (see Fig. 2 of Kamieneski et al. 2023), which is the reason why the only visible image of clump has been associated to the bulge in source plane. Consequently, we caution on using solely photometric signatures (shapes or colors) to draw conclusions on the lensing origin of features in the arcs. The use of incorrect pairs of multiple images is likely to bias the inference on the source morphology and mass model parameters. However, depending on the application and the required precision of the measurements, these biases may not always affect the conclusions of the analyses. In particular, the analysis of Kamieneski et al. (2023) is likely unaffected, as it focuses on large-scale properties of the source, such as the stellar mass, star formation rate and color gradients over kiloparsec scales.
@@ 7 Summary and conclusion
In this work, we model in details the prominent strong lens system El Anzuelo in the vicinity of the massive cluster El Gordo. We do so using high-resolution and multi-band JWST/NIRCam imaging data obtained from the PEARLS program (Windhorst et al. 2023) and MUSE/VLT specstroscopic data (Caminha et al. 2023). This system is composed of three elliptical galaxies as main deflectors (L1, L2, L3) at redshift $z_{\rm d}\approx 0.9$ and a strongly lensed dusty star forming galaxy at redshift $z_{\rm s}\approx 2.3$. We summarize below the novelty of our analysis and its main results:
@@
  • Our high-resolution multi-band source model allows us to detect and locate a small luminous clump, with a maxmimal size of $400$ pc at the source redshift. To our knowledge, this may be one of the first clumps of this size located in a dusty galaxy at redshift $z\sim 2.3$. Only one out of the four multiple images of the clump is visible in the data, but its detection is possible thanks to extended lensing constraints, color information and high resolution of JWST/NIRCam.
@@ Acknowledgements.
AG acknowledges funding and support by the Swiss National Science Foundation (SNSF). AG thanks Philipp Frank and Gordian Edenhofer for fruitful discussions and help with NIFTy. AG also thanks Patrick Kamieneski for providing additional details on their lens model of El Anzuelo, as well as Rogier Windhorst, Jake Summers, and Anton Koekemoer from the PEARLS team for providing additional details on reduced JWST data products. JK and SHS acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC 2094 – 390783311. JR acknowledges financial support by the German Federal Ministry of Education and Research (BMBF) under grant 05A20W01 (Verbundprojekt D-MeerKAT). JK also acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101071643). SHS thanks the Max Planck Society for support through the Max Planck Fellowship. This research has also made use of SciPy (Virtanen et al. 2020), NumPy (Oliphant 2006; Van Der Walt et al. 2011), Matplotlib (Hunter 2007), Astropy (Astropy Collaboration et al. 2013; Price-Whelan et al. 2018), Jupyter (Kluyver et al. 2016) and GetDist (Lewis 2019).
Appendix A Complementary results and discussion
A.1 Surface brightness properties of the deflectors

### v2

@@ 5.4 Lensed image families and clumps in the source
We are able to confirm that parts of the background galaxy are multiply imaged five times (excluding central images); in Fig. 9 this is shown by image families 1 and 3. This peculiar lensing configuration has been first suspected in Kamieneski et al. (2023). We check if these fifth images are magnified by computing their model-predicted magnification from our fiducial mass model. We obtain magnifications of $\mu\approx-3$ for image 1d and $\mu\approx-5$ for image 3b, thus we find there are magnified. We note that this “exotic” lensing configuration typically arises in systems with multiple deflectors and has been the topic of several theoretical works in the past (e.g., Shin & Evans 2008; Orban de Xivry & Marshall 2009; Bozza et al. 2020). We briefly discuss in Appendix C.4 potential implications regarding this lensing configuration.
5.5 Detection of multiply imaged clumps in the source
@@
We find that one of the images used in Kamieneski et al. (2023) as being part of the image family 1 (their image 1b) in fact does not belong to this set of multiple image according to our lens model. On the one hand, we find that their image 1d is in fact composed of the two images denoted by 1a and 1c, which appear like merging while crossing the tangential critical line. On the other hand, their image 1b corresponds to our image 4a in Fig. 9, and thus originates from a distinct compact bright clump in the source, located at the north of the central bulge. We note that this clump has similar colors as the bulge of the source galaxy, which is why it could have been incorrectly attributed to a lensed image of the bulge. Interestingly, not all images of this clump are visible in the data. We believe it is mostly due to a combination of a the larger FWHM of the PSF in redder bands, and stronger contamination by the deflectors light. To better visualize this, we show in Fig. 12 a comparison between our the lensed and convolved source model and the corresponding model before convolution by the PSF (in F356W although it is similar in other filters). We see that only image 4a remains clearly visible after convolution. Moreover, the lensed images appear either close to the bright bulge of the source (image 4d), or close to regions heavily contaminated by the deflectors light (images 4b and 4c); therefore, they are not seen in the data. From one of our higher resolution source models (with $n_{\rm src}=300$), we estimate that the physical diameter (in projection) of the clump is approximately $400$ pc at redshift $z_{\rm s}=2.291$. We measure that this clump is located at approximately $1.2$ kpc from the source centroid. While the size estimate is an upper-bound to the true size—since increasing the source resolution tends to make it more compact—this clump in the source is remarkably small. It is comparable in size to some of the structures in the jets of lensed sources analyzed in Powell et al. (2022) using VLBI observations with milli-arcsecond resolution (against 30 milli-arcsecond resolution with JWST/NIRCam data).
@@
Finally, we note that our extended source model can be used to extract the SED of compact unresolved structures in the arc. We show in Appendix D the case of one of the brightest point-like feature in El Anzuelo (image 2b in Fig. 9). As our pixelated source model, at its fiducial resolution ($n_{\rm src}=100$), does not capture unresolved features by construction, it can be directly subtracted from the data to separate the extended source component from these unresolved features. It is then easier to perform photometric measurements and extract the SED of such compact source regions, which are important tracers of the star formation history of the galaxy (as shown in Kamieneski et al. 2023).
@@ 5.7 Release of lens models within the COOLEST standard
Previous sections describe our modeling and inference approach with a level of details that aim at maximizing the reproducibility of our analysis. With this idea in mind, we also publicly release all lens models [^13] https://github.com/aymgal/ElAnzuelo_modeling_public following the strong lensing standard COOLEST (Galan et al. 2023). These models can then be easily loaded using the Python package coolest [^14] https://github.com/aymgal/COOLEST for further analyses, in particular with any modeling codes that share an interface with COOLEST (see the documentation for a list). We encourage future lens models to be released under this standard, in order to facilitate comparison with the present analysis.
@@ 6.1 Mass density profile of elliptical galaxies
The three foreground deflectors if El Anzuelo have typical photometric signatures of elliptical galaxies and have been confirmed to be members of the massive cluster El Gordo (e.g., Caminha et al. 2023). In this work, we model the mass distribution of these galaxies using elliptical power-law profiles perturbed by the cluster mass distribution and an additional external shear. As presented in Sect. 5 (Table 2), we infer logarithmic density slopes for L1 and L2 of $\gamma_{\rm L1}=2.23\pm 0.05$ and $\gamma_{\rm L2}=2.21\pm 0.04$, respectively. These values are inconsistent with an isothermal slope, and $\gamma_{\rm L2}$ is $2.7\sigma$ away from the mean value $\langle\gamma\rangle=2.075\pm 0.024$ of Etherington et al. (2023) inferred from a subset of SLACS (Bolton et al. 2006) and GALLERY (Shu et al. 2016) systems. As detailed in Sect. 4, we infer our uncertainties by marginalizing over an ensemble of model variations with Bayesian weighting. We also looked for potential systematic effects and did not find statistically significant changes in the posteriors distributions by running additional models. Here, we first compare the precision of these measurements with previous analyses of similar imaging data, then discuss the possible origins of these results in terms of galaxy structure.
@@
That elliptical galaxies have nearly isothermal radial density profiles in their inner region has been observed in isolated deflectors, such as those from the SLACS (Bolton et al. 2006) and SL2S (Gavazzi et al. 2012) samples. The three main deflectors of El Anzuelo are, however, not isolated but genuine members of El Gordo that are gravitationally interacting. Therefore, it may not be surprising to measure density slopes that deviate from an isothermal profile. For most of isolated lens systems from SLACS (or comparable samples) we note that the ratio $\bar{r}_{\rm E/eff}$ between the Einstein radius and the half-light radius, $\bar{r}_{\rm E/eff}\equiv\theta_{\rm E}/\theta_{\rm eff}\lesssim 1$, meaning that lensing overall constrains the inner region of the lens galaxies, within their half-light radius (e.g., see the middle panel of Fig. 15 from Etherington et al. 2022, for SLACS and GALLERY samples). For the case of El Anzuelo, L1 and L2 both have $\bar{r}_{\rm E/eff,L1}\approx 1.6$ and $\bar{r}_{\rm E/eff,L2}\approx 1.3$ respectively (as visualized in Fig. 11), meaning that lensing provides constraints at overall larger radii compared to lens galaxies from the aforementioned samples [^16] The effective radius for L2 may be underestimated due to due blending with the lensed arcs (see Appendix C.1). However, due to the lensing boost by the cluster, the ratio $\bar{r}_{\rm E/eff,\rm L2}$ is likely larger than unity.. The main reason for El Anzuelo deflectors to have $\bar{r}>1$ is that their apparent lensing effect is boosted by the mass density of the host cluster, leading to the appearance of an Einstein ring that is larger than if L1 or L2 were isolated galaxies. This effect can be qualitatively confirmed by noting the difference in $\theta_{\rm E}$ values between the “Shear only” and “Cluster+Shear” models (e.g., Fig. 10): including the cluster mass distribution in the model significantly reduces $\theta_{\rm E}$ values for L1 and L2 in order to fit the imaging data. Therefore, the observed lensing features probe the mass distribution of the deflectors at sensibly larger radii, compared to isolated deflectors.
@@
6.2 Approximating the host cluster with an external shear
Figure 13: Shear orientation and strength over the field of view of El Anzuelo. For the “Shear only” model, the shear field is a uniform external shear with strength $\gamma_{\rm ext}$; for the “Cluster only”, it is the shear field from the cluster model of Caminha et al. (2023); for the “Cluster+Shear” model, it is the combined effect from the uniform external shear and the cluster shear. For the latter two models, the averaged shear strength $\langle\gamma_{\rm shear}\rangle$ is also indicated in the legend. The absolute length of the shear markers is arbitrary, but their relative size is respected. The blue ellipse indicates the orientation of the mass density of L1. On the left of the plot, we show the directions pointing towards the two main dark matter halo component of the El Gordo cluster. Note that the shear orientation is consistent with the position of the cluster scale components.
@@ 6.3 Using extended arcs instead of point-like constraints
In Sect. 5.4 we compared in more details our results with the lens model of Kamieneski et al. (2023), especially for three image families. In their analysis, the authors used the 8 NIRCam filters to look for compact features in image plane that have consistent colors, and considered those as multiple images of same region in the source. For their image family 1, they used a simplified initial lens model to confirm the five-image configuration, interpreted as a confirmation of the origin of images 1a to 1e being lensed versions of the bright central region of the source galaxy. In Sect. 5.5, we show that image 1b does not originate from the center of the source, but rather from an off-centered compact clump of size $\lesssim 400$ pc. In NIRCam images, this clump has similar colors as the central bulge of the source (see Fig. 2 of Kamieneski et al. 2023), which is the reason why the only visible image of the clump has been associated to the bulge in source plane. Consequently, we caution on using solely photometric signatures (shapes or colors) to draw conclusions on the lensing origin of features in the arcs. The use of incorrect pairs of multiple images is likely to bias the inference on the source morphology and mass model parameters. However, depending on the application and the required precision of the measurements, these biases may not always affect the conclusions of the analyses. In particular, the analysis of Kamieneski et al. (2023) is likely unaffected, as it focuses on large-scale properties of the source, such as the stellar mass, star formation rate and color gradients over kiloparsec scales.
Our analysis makes use of all pixels that contain lensing features in the six NIRCam filters we selected. Another approach, which can be seen as intermediate between point-like and pixel-level modeling, is to locally extract shape information from individual multiple images to constrain local properties of the lens potential. This idea has been formalized and explored in several works (Wagner & Bartelmann 2016; Tessore 2017; Fleury et al. 2019; Birrer 2021), although with limited applications to real observations so far (e.g., Wagner et al. 2018). The main advantage in locally constraining the lens potential is to release most of the assumptions regarding the global lens mass distribution, at the cost of additional assumptions regarding the symmetry and compactness of the source (or individual features within it). In systems such as El Anzuelo, the lensed source is likely too asymmetric to directly apply such a formalism. However, it may be possible to combine constraints from extended source modeling with macro-model independent corrections based on the shape of individual lensed resolved clumps within the arc (see e.g., Lin et al. 2023, for an automated multi-band extraction of lensed features).
@@ 7 Summary and conclusion
In this work, we model in details the prominent strong lens system El Anzuelo in the vicinity of the massive galaxy cluster El Gordo. We do so using high-resolution and multi-band JWST/NIRCam imaging data obtained from the PEARLS program (Windhorst et al. 2023) and MUSE/VLT specstroscopic data (Caminha et al. 2023). This system is composed of three elliptical galaxies as main deflectors (L1, L2, L3) at redshift $z_{\rm d}\approx 0.9$ and a strongly lensed dusty star-forming galaxy at redshift $z_{\rm s}\approx 2.3$. We summarize below the novelty of our analysis and its main results:
@@
  • Our high-resolution multi-band source model allows us to detect and locate a small luminous clump located at $1.2$ kpc from the source centroid and with a maxmimal diameter of $400$ pc at the source redshift. To our knowledge, this may be one of the first clumps of this size located in a dusty galaxy at redshift $z\sim 2.3$. Only one out of the four multiple images of the clump is visible in the data, but its detection is possible thanks to extended lensing constraints, color information and the high resolution of JWST/NIRCam.
@@ Acknowledgements.
The authors warmly thank the anonymous referee for the many useful suggestions that improved this manuscript. AG acknowledges funding and support by the Swiss National Science Foundation (SNSF). AG thanks Philipp Frank and Gordian Edenhofer for fruitful discussions and help with NIFTy, as well as Claudio Grillo for his thoughts on the results. AG also thanks Patrick Kamieneski for providing additional details on their lens model of El Anzuelo, as well as Rogier Windhorst, Jake Summers, and Anton Koekemoer from the PEARLS team for providing additional details on reduced JWST data products. AG thanks the Max Planck Computing and Data Facility for support. JK and SHS acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC 2094 – 390783311. JR acknowledges financial support by the German Federal Ministry of Education and Research (BMBF) under grant 05A20W01 (Verbundprojekt D-MeerKAT). JK also acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101071643). SHS thanks the Max Planck Society for support through the Max Planck Fellowship. This research has also made use of SciPy (Virtanen et al. 2020), NumPy (Oliphant 2006; Van Der Walt et al. 2011), Matplotlib (Hunter 2007), Astropy (Astropy Collaboration et al. 2013; Price-Whelan et al. 2018), Jupyter (Kluyver et al. 2016) and GetDist (Lewis 2019).
Appendix A Weighting used in color composite images
In Figs. 1 and 7, we constructed color composite images of JWST and HST datasets. The arbitrary weighting of individual bands, chosen for visualization purposes only, is the following:
[…]
The Lupton et al. (2004) scheme is then used to create the final color images.
Appendix B Uncertainty about the source redshift
As mentioned in Sect. 3.1, we assume a source redshift of $z_{\rm s}=2.291$ throughout this work. However, Kamieneski et al. (2023) also indicate that an alternative redshift value, although less likely, can be $z_{\rm s}=3.388$ if the emission line detected in ALMA data corresponds to the CO(4-3) transition instead of CO(2-1). Here we discuss what would be the changes implicated by this alternative redshift.
Assuming a different source redshift impacts the Einstein radii estimates obtained from stellar velocity dispersion measurements (through the angular diameter distance $D_{\rm s}$, see Eq. 2). Assuming $z_{\rm s}=3.388$, and a realistic contribution from the cluster to the total potential would increase the Einstein radii to $\theta_{\rm E,L1}=1\aas@@fstack{\prime\prime}25\pm 0\aas@@fstack{\prime\prime}05$ and $\theta_{\rm E,L3}=0\aas@@fstack{\prime\prime}58\pm 0\aas@@fstack{\prime\prime}06$ for L1 and L3, respectively. Compared to the values from Table 1, this corresponds to an increase of $\sim 24\%$ for L3. The disagreement between the lensing and stellar kinematics estimates of $\theta_{\rm E,L3}$ discussed in Sect. 5.3 would increase to $3.7\sigma$.
The second change caused by a different source redshift relates to the physical sizes of the reconstructed lensed features in source plane. If one assumes $z_{\rm s}=3.388$, the physical size of the clump detected in Sect. 5.5 and its distance to the source centroid would reduce to $\sim 360$ pc and $\sim 1.1$ kpc, respectively. In other words, the alternative source redshift would imply that the clump is smaller and closer to the galaxy center.
The last consequence of the uncertain source redshift is the rest-frame wavelength quoted in Fig. 17, which shows the SED of an unresolved source feature. Since we do not specifically use the SED, this has no impact on our results.
Appendix C Complementary results and discussion
C.1 Surface brightness properties of the deflectors